# RSRNav: Reasoning Spatial Relationship for Image-Goal Navigation

### v1

@@ 4.3 Ablation Study
Analysis of Dense Correlation.    Evolution from “Minimalist Correlation” to “Dense Correlation”, there are two key points in the construction of dense correlation: cross-correlation and fine-grained correlation. Here, we explore the contribution of the two key points to performance improvement. As shown in Table 5, both can bring performance gains, fine-grained correlation significantly improves information content while cross correlation allows the global relation between goal and egocentric images. Although they can improve performance separately, together, they can maximize benefits. Essentially, both are increasing the amount of correlation information, demonstrating that as the correlation information gets richer, the performance gets better.
@@
Analysis of Direction-aware Correlation.    The first and last rows of Table 6 present the performance with “Dense Correlation” and “Direction-aware Correlation”, respectively. The results of the first three lines show that when we simply replace cross-correlation with the LookUp operation, it brings about performance degradation, which only works if we increase the size of the feature map. We then conduct experiments on the correlation pyramid as shown in the last three rows of Table 6. Shows that as the number of layers increases, the navigation performance improves. As seen in Fig 3, the multi-layer pyramid yields a multi-scale directional correlation; the more layers there are, the richer the correlation information is. These demonstrate the validity of our direction-aware correlation, and a more qualitative analysis of direction information is in the visualizations below.
Fig 4: Compare the number of parameters, as well as the SPL on Gibson.
@@
Performance vs. Model Size.    Smaller AI models bring significant benefits to real-world deployments by requiring less computational power, enabling them to run on a wider range of devices with less energy consumption. This is even more important for an embodied intelligence that requires visual navigation, as it has to carry other operational facilities in addition to what it needs to navigate, so the computational resources required for navigation are stringent. As shown in Fig 4, we compare the number of parameters, as well as the SPL on Gibson for [5, 23, 4, 6, 7]. Our RSRNav outperforms existing methods and has the smallest number of parameters among all methods, showing potential for real-world applications.
@@
More Extreme “user-matched goal” Setting.    When used in the real world, there should not be too much restriction on the goal camera as the user can shoot the goal image at any height and from any angle. In addition to the two goal settings in the main text, i.e. “agent-matched goal” and “user-matched goal”, we add a more extreme setting of goal camera in this part. Specifically, we sample the height of the goal camera from $\mathcal{U}(0.8m,1.5m)$, the pitch delta from $\mathcal{U}(-45^{\circ},45^{\circ})$, and the HFOV from $\mathcal{U}(60^{\circ},120^{\circ})$. The results reported in Table 7 show that our method significantly surpasses both the baseline and state-of-the-art method FGPropm in performance.
Table 7: Additional comparison under “extreme setting” on Gibson.
@@
Fig 5: Comparison of convergence rates and tests results on metrics “Distance to Goal” and “Angle to Goal”. The former metric measures the distance between the agent and the goal when stopping; the latter metric is the angle between the agent pose and the goal view when stopping. We also provide the average episode test results, plotted to the right of the convergence curve.
4.4 Analysis of Correlation-based method for ImageNav
We analyze human-intuitive navigation, and inspired by it we identify two limitations of previous methods: (1) Semantic feature vectors inadequately convey valid direction information for navigation, resulting in inefficient and superfluous actions; (2) Performance decreases dramatically when encountering viewpoint inconsistencies caused by differences in camera settings between training and application. Previous methods have not explicitly model the comparison between the current observation to the goal. They only provide the feature vectors, putting the full burden of comparison on the policy network and complicating the training. Semantic feature vectors lack precise spatial details, resulting in unclear directions for the policy and potentially causing unnecessary actions during navigation towards the goal. In this part, we analyze how our proposed correlation-based method addresses these challenges. For the first challenge, we analyze and experimentally verify that the correlation is the key information to guide navigation actions. Therefore, instead of directly passing the features in the perception step, we construct the correlation between them and then allow the policy network to predict action based on it. This can significantly reduce the learning burden of the policy network. As shown in Fig 5, “Navigation Via Dense Correlation” and “Navigation Via Direction-aware Correlation” converge significantly faster with sufficient correlation information given to the policy network. For the second challenge, whether the angle can be adjusted accurately according to the goal can be reflected most intuitively in the “Angle to Goal” metric, which is the angle between the agent pose and the goal view when stopping. Comparison of convergence curves and test results for this metric, as shown in Fig 5, demonstrates that correlation can help the agent adjust the pose accurately. Then, after we introduce direction-aware correlation, the precision of the agent’s pose becomes higher.
Fig 6: Comparison of Visualization Results. The gray area in the scene represents the area scanned by the egocentric camera while the agent is traveling. The agent’s navigational path is indicated by a blue line that gets brighter as the time steps get longer (from dark blue to light blue). These are the results under the “user-matched camera” setting.
Fig 7: Comparison of the three methods, with the target viewpoint image on the left, the image captured by agent’s own camera in the center, and the topdown image on the right.
4.5 Visualization
We visualize the navigation results under the “user-matched goal” setting, with the navigation path visualized on the top-down maps. We compare our RSRNav with the other two methods for different levels of difficulty in navigation in Fig 6 (from right to left, navigation goes from easy to hard).
@@ 4.5 Visualization
We intercepted two of these navigation result comparisons as shown in Fig 7. The success of a navigation task is judged by the distance to the target after the “stop” command. However, as shown in the red box in the figure, even though the previous method can successfully navigate to the target location, it is still not precise enough for the angle of view, whereas we can make the navigation angle more precise, thanks to the modeling of the orientation information.
5 Conclusion
@@
Acknowledgments
This work was supported in part by National Science and Technology Major Project under Grant 2024YFB4708100, National Natural Science Foundation of China under Grants 62088102, 12326608 and 62106192, Natural Science Foundation of Shaanxi Province under Grant 2022JC-41, Key R&D Program of Shaanxi Province under Grant 2024PT-ZCK-80, and Fundamental Research Funds for the Central Universities under Grant XTR042021005.

### v2

@@ IV-C Ablation Study
Analysis of Dense Correlation.   Evolution from “Minimalist Correlation” to “Dense Correlation”, there are two key points in the construction of dense correlation: cross-correlation and fine-grained correlation. Here, we explore the contribution of the two key points to performance improvement. As shown in Table V, both can bring performance gains, fine-grained correlation significantly improves information content while cross correlation allows the global relation between goal and egocentric images. Although they can improve performance separately, together, they can maximize benefits. Essentially, both are increasing the amount of correlation information, demonstrating that as the correlation information gets richer, the performance gets better.
Figure 5: Compare the number of parameters, as well as the SPL on Gibson.
@@
Analysis of Direction-aware Correlation.   The first and last rows of Table VI present the performance with “Dense Correlation” and “Direction-aware Correlation”, respectively. The results of the first three lines show that when we simply replace cross-correlation with the LookUp operation, it brings about performance degradation, which only works if we increase the size of the feature map. We then conduct experiments on the correlation pyramid as shown in the last three rows of Table VI. Shows that as the number of layers increases, the navigation performance improves. As seen in Figure 3, the multi-layer pyramid yields a multi-scale directional correlation; the more layers there are, the richer the correlation information is. These demonstrate the validity of our direction-aware correlation, and a more qualitative analysis of direction information is in the visualizations below.
Figure 6: Comparison of Visualization Results. The gray area in the scene represents the area scanned by the egocentric camera while the agent is traveling. The agent’s navigational path is indicated by a blue line that gets brighter as the time steps get longer (from dark blue to light blue). These are the results under the “user-matched camera” setting.
Figure 7: Comparison of three methods. The left side shows the target viewpoint image, the middle shows the image captured by the agent’s own camera after reaching the target location, and the right side shows the overhead map. As can be seen from the red boxes, although FGPrompt found the target location and was marked as successful, there is a deviation in accuracy.
@@
Performance vs. Model Size.   Smaller AI models bring significant benefits to real-world deployments by requiring less computational power, enabling them to run on a wider range of devices with less energy consumption. This is even more important for an embodied intelligence that requires visual navigation, as it has to carry other operational facilities in addition to what it needs to navigate, so the computational resources required for navigation are stringent. As shown in Figure 5, we compare the number of parameters, as well as the SPL on Gibson for [7, 29, 6, 8, 9]. Our RSRNav outperforms existing methods and has the smallest number of parameters among all methods, showing potential for real-world applications.
@@
More Extreme “user-matched goal” Setting.   When used in the real world, there should not be too much restriction on the goal camera as the user can shoot the goal image at any height and from any angle. In addition to the two goal settings in the main text, i.e. “agent-matched goal” and “user-matched goal”, we add a more extreme setting of goal camera in this part. Specifically, we sample the height of the goal camera from $\mathcal{U}(0.8m,1.5m)$, the pitch delta from $\mathcal{U}(-45^{\circ},45^{\circ})$, and the HFOV from $\mathcal{U}(60^{\circ},120^{\circ})$. The results reported in Table VII show that our method significantly surpasses both the baseline and state-of-the-art method FGPropm in performance.
TABLE VII: Additional comparison under “extreme setting” on Gibson.
@@
IV-D Analysis of Correlation-based method for ImageNav
We analyze human-intuitive navigation, and inspired by it we identify two limitations of previous methods: (1) Semantic feature vectors inadequately convey valid direction information for navigation, resulting in inefficient and superfluous actions; (2) Performance decreases dramatically when encountering viewpoint inconsistencies caused by differences in camera settings between training and application. FGPrompt has not explicitly modeled the comparison between the current observation to the goal. They only provide the feature vectors, putting the full burden of comparison on the policy network and complicating the training. Semantic feature vectors lack precise spatial details, resulting in unclear directions for the policy and potentially causing unnecessary actions during navigation towards the goal. In this part, we analyze how our proposed correlation-based method addresses these challenges. For the first challenge, we analyze and experimentally verify that the correlation is the key information to guide navigation actions. Therefore, instead of directly passing the features in the perception step, we construct the correlation between them and then allow the policy network to predict action based on it. This can significantly reduce the learning burden of the policy network. As shown in Figure 4, “Navigation Via Dense Correlation” and “Navigation Via Direction-aware Correlation” converge significantly faster with sufficient correlation information given to the policy network. For the second challenge, whether the angle can be adjusted accurately according to the goal can be reflected most intuitively in the “Angle to Goal” metric, which is the angle between the agent pose and the goal view when stopping. Comparison of convergence curves and test results for this metric, as shown in Figure 4, demonstrates that correlation can help the agent adjust the pose accurately. Then, after we introduce direction-aware correlation, the precision of the agent’s pose becomes higher.
IV-E Visualization
We visualize the navigation results under the “user-matched goal” setting, with the navigation path visualized on the top-down maps. We compare our RSRNav with the other two methods for different levels of difficulty in navigation in Figure 6 (from right to left, navigation goes from easy to hard).
@@ IV-E Visualization
We intercepted two of these navigation result comparisons as shown in Figure 7. The success of a navigation task is judged by the distance to the target after the “stop” command. However, as shown in the red box in the figure, even though the previous method can successfully navigate to the target location, it is still not precise enough for the angle of view, whereas we can make the navigation angle more precise, thanks to the modeling of the orientation information.
V Conclusion